A skater glides along a circular path of radius 2.50m . If she coasts around one half of the circle, find the distance in m from the starting location


1
Expert's answer
2022-04-11T12:07:39-0400

The distance traveled by a skater from the starting location

"d=2\pi R/2=\pi R\\
=3.14*2.5=7.85\:\rm m"